in a survey of women in a certain country ages (20-29) the mean height was 62.9 inches with a standard deviation of 2.76. Answer the following questions about the specified normal distribution.

a. what height represents the 95th percentile?
b. what height represents the first quartile?
Answer #1

Solution:- Given that mean = 62.9inches, standard deviation = 2.76

X = μ + (Z*σ)

a) z = 1.645

X = 62.9 + (1.645*2.76) = 67.4402

b) z = -0.6745

X = 62.9 + (-0.6745*2.76) = 61.0384